### v1

# Looking Beyond Visible Cues: Implicit Video Question Answering via Dual-Clue Reasoning

Abstract
Video Question Answering (VideoQA) aims to answer natural language questions based on the given video, with prior work primarily focusing on identifying the duration of relevant segments, referred to as explicit visual evidence. However, explicit visual evidence is not always directly available, particularly when questions target symbolic meanings or deeper intentions, leading to significant performance degradation. To fill this gap, we introduce a novel task and dataset, Implicit Video Question Answering (I-VQA), which focuses on answering questions in scenarios where explicit visual evidence is inaccessible. Given an implicit question and its corresponding video, I-VQA requires answering based on the contextual visual cues present within the video. To tackle I-VQA, we propose a novel reasoning framework, IRM (Implicit Reasoning Model), incorporating dual-stream modeling of contextual actions and intent clues as implicit reasoning chains. IRM comprises the Action-Intent Module (AIM) and the Visual Enhancement Module (VEM). AIM deduces and preserves question-related dual clues by generating clue candidates and performing relation deduction. VEM enhances contextual visual representation by leveraging key contextual clues. Extensive experiments validate the effectiveness of our IRM in I-VQA tasks, outperforming GPT-4o, OpenAI-o3, and fine-tuned VideoChat2 by 0.76%, 1.37%, and 4.87%, respectively. Additionally, IRM performs SOTA on similar implicit advertisement understanding and future prediction in traffic-VQA. Datasets and codes are available for double-blind review in anonymous repo: https://github.com/tychen-SJTU/Implicit-VideoQA.
1 Introduction
Video Question Answering (VideoQA) is designed to comprehend visual content and generate accurate answers by reasoning over the provided video. To facilitate reasoning through more relevant visual segments, an increasing number of VideoQA works have focused on locating key visual segments, referred to as explicit visual evidence. With the explicit evidence, reasoning models (e.g., Large Language Models) can derive more accurate answers sevila ; gqa1 ; relation . Furthermore, recent works have introduced Grounded Video Question Answering (Grounded-VQA) task nextgqa ; lita ; etbench ; rextime ; cgbench , which employs supervised training to locate explicit visual evidence more precisely.
Figure 1: Examples of Implicit VQA. Implicit VQA is more challenging due to the non-existence of explicit visual evidence. This happens when questions target symbolic meanings (left figure) or deeper intentions (right figure).
However, the performance of VideoQA is significantly compromised when explicit visual evidence is difficult to locate directly. This challenge arises from several factors, including but not limited to ambiguous visual segments ambiguous , perspective shifts perspective , and occlusions occlusions . Moreover, grounding explicit evidence becomes particularly challenging for questions that involve symbolic meanings or underlying intentions auvid ; mmau ; vvtut . In these cases, the direct identification of explicit evidence may be confounded by spurious relationships, underscoring the need for a deeper reasoning approach that goes beyond surface-level connections.
As shown in Fig. 1, the first scenario illustrates a case where direct visual grounding faces challenges, as there are no direct reasoning pathways linking the intent of “go to McDonald’s” to visual segments. Instead, through deeper reasoning, we can uncover indirect and implicit relations: the similar shape between the McDonald’s logo and the double curled eyelashes. The second scenario further emphasizes the importance of deeper contextual analysis, where the advertisement’s intent is conveyed through implicit environmental cues that evoke warmth and comfort.
To address these challenges, we set up a new task: Implict-Video Question Answring (I-VQA), which aims to infer an appropriate answer to implicit video questions, where explicit visual evidence corresponding to the given question cannot be directly accessed or located. Meanwhile, we contribute a new dataset for the training and evaluation of I-VQA by reorganizing videos from the Grounded-VQA datasets and removing the annotated visual evidence segments. However, to our knowledge, no existing solutions can directly tackle general implicit questions.
To this end, we draw inspiration from works in similar evidence-insufficient task: visual commonsense reasoning. These approaches enhance reasoning by leveraging contextual visual information and commonsense to infer additional relevant details visualcomet ; video2commonsense ; visualcommon2 . Moreover, in the video modality, richer temporal contextual inputs make it even more crucial to effectively incorporate textual information. When further considering the utilized form, cognitive science suggests Cognitive that higher-level information has greater potential for guiding reasoning. As shown in  Fig. 2, the intention conveyed in the preceding context is often reflected in subsequent actions. We can infer implicit actions by extracting intention from the preceding context, while also validating the coherence of these implicit intentions through actions in the following context. Therefore, we argue that introducing (1) context action involvement clues (2) context underlying intention clues as additional reasoning pathways is critical for implicit reasoning in the I-VQA task.
Based on this intuition, we introduce IRM, (Implicit Reasoning Model), which employs dual-stream modeling of contextual actions and intent clues as implicit reasoning chains to capture deeper underlying semantics. Specifically, we propose the Action-Intent Module (AIM), which generates textual clues representing explicit contextual actions and their underlying intentions. However, these clue candidates may be affected by factors such as hallucination MECD ; hallucination and language bias biasgpt ; biasgpt2 . To mitigate error propagation and reduce potential misleading tendencies, we incorporate a visual information verification module within AIM. This module refines dual-clue candidates by verifying them against visual information, leveraging global temporal perception for more accurate validation.
Furthermore, the relationships among sub-events in a video are diverse, leading to variations in the importance of context clues. Therefore, we propose allocating greater attention to the most relevant context clues. Specifically, we introduce a relationship classifier within AIM, which evaluates the relevance of context clues to the implicit question. Only those clues deemed beneficial for implicit question answering are retained, ensuring more effective reasoning.
Figure 2: Dual Clues Aid in Answering Implicit Video Questions. When addressing implicit problems, direct reasoning frequently fails (as indicated by the red arrows). In contrast, our proposed context clue-aided reasoning method (as indicated by the green arrows) first infers the underlying intent from the preceding context, and then verifies this intent using subsequent actions.
In addition to constructing textual clues, we introduce the Visual Enhancement Module (VEM) to enhance the visual dual-clue representation. VEM applies attention mechanisms between the originally extracted contextual visual information and the refined clues. The LLM then performs reasoning with the augmented dual-clue and enhanced visual information. Notably, the enhanced visual information can be selectively fed back into the AIM, further mitigating hallucinations and improving relationship deduction. This iterative process strengthens the overall modeling of clues.
Extensive experiments validate the effectiveness and generalizability of IRM for implicit reasoning. IRM performs SOTA in both the proposed I-VQA and current advertisement understanding datasets. Additionally, the generalization test highlights the strong performance of IRM in explicit VideoQA, particularly in predicting outcomes in SUTD-Traffic and the corresponding MVBench subset.
2 Related Work
Video Question Answering. Video question answering, as a key representative of video reasoning tasks, has seen rapid development. Recent advancements in VideoQA have been driven by the integration of pre-trained visual encoders with large language models videollama ; stllm ; vqa_1 ; vqa_2 ; vqa_3 . Notably, several works sevila ; gqa1 ; relation have focused on identifying key visual segments within the temporal domain during the reasoning process, enabling more accurate reasoning. The Grounded Video Question Answering (Grounded-VQA) task nextgqa ; lita ; etbench ; rextime ; cgbench ; timecraft ; vqa_4 ; video_thought further strengthens this approach by incorporating human annotations for supervised training. However, the performance of these methods can significantly degrade when explicit evidence is difficult to locate directly. Therefore, we propose the I-VQA task, which is specifically designed to handle the challenging implicit VideoQA.
Visual Context Reasoning. Visual context reasoning refers to scenarios where the available information is insufficient for supporting the visual understanding task. Current approaches to visual context reasoning primarily focus on leveraging commonsense knowledge to infer additional visible visual information visualcomet ; video2commonsense ; visualcommon ; visualcommon2 . However, these methods are prone to reporting bias and relying on commonsense assumptions can sometimes lead to misguided conclusions. Another approach models the relationships between contexts, using known context information to infer masked events in video captioning VAR or video causal relation reasoning MECD ; mecd+ . However, these methods are susceptible to error accumulation, which can negatively affect subsequent reasoning. We offer a more robust solution, rather than explicitly inferring additional visual data, we model informative contextual clues related to action and intention. These clues form reasoning chains that capture deeper underlying semantics, while error accumulation is mitigated through a rigorous refinement process.
Implicit Video Reasoning. The reasoning tasks related to implicitness in videos mainly focus on inferring implicit attributes of objects, such as mass and charge in virtual videos. ComPhy comphy requires models to identify intrinsic properties and make dynamic predictions based on these. CRIPP-VQA cripp requires physical properties to be learned from video rather than explicitly expressed in the question. ACQUIRED ACQUIRED is a combination of the two forms mentioned above. Another similar task is the PSAV adsdataset , which introduces the persuasion strategy identification task in advertisement understanding, the model needs to distinguish the implicit strategy of the advertisement video among the strategies defined by Persuasion . However, our I-VQA focuses on more general implicit questions regarding daily-life videos. Moreover, our task eliminates the possibility of taking explicit evidence as shortcuts, ensuring the rigor of implicit reasoning.
3 I-VQA Task
In Sec. 3.1, we introduce our I-VQA task setting, while in Sec. 3.2 we present the data sources, the principles of collection and annotation, and the characteristics of our dataset.
3.1 Task Definition
We define the task of I-VQA as: taking a video without explicit visual evidence $v~{}=~{}v_{0}\setminus v_{e}$, a question $q$, and a corresponding answer set $\mathbb{A}$ as input, the model is required to deduce the correct answer ${a}^{\ast}$ from the answer set $\mathbb{A}$. Where $v_{0}$ indicates the complete video with a duration of $t$, and $v_{e}$ indicates the explicit visual evidence that occurs during $\mathbb{T_{e}}=\{t_{e}^{1},...,t_{e}^{n}\},where~{}t_{e}^{i}=[t_{0},t_{1}],0\leq t_{0}<t_{1}\leq t$. The reasoning process can be represented by: $a^{*}=\arg\max_{a\in\mathbb{A}}~{}f_{w}(a|q,v,\mathbb{A})$, where $f_{w}$ represents a mapping function with learnable parameters $w$.
Moreover, to ensure the implicitness, we appropriately extend the explicit visual evidence, where the certain excluded timestamp $t_{e}^{i}$ can be represented by: $t_{e}^{i}=[max(0,t_{0}-{T}/{\sigma}),min(t,t_{1}+{T}/{\sigma})]$, where $\sigma$ indicates a constant value. Besides, following previous VideoQA tasks videollava ; qwen ; pllava ; stllm , we introduce the open-ended evaluation for a subset in the I-VQA task. After an extremely strict manual assessment, the open-ended evaluation subset is identified, which satisfies the condition of a unique correct answer without any imposed restrictions. The evaluation is represented by: $a_{o}=f_{w}(q,v)$. $a_{o}$ is the generated responses, which can be in free-form language for open-ended VideoQA.
3.2 Task Dataset
Data Source. The I-VQA task involves question-answer pairs where key visual evidence is invisible. To set up the task setting for implicit questions, we systematically reorganize and leverage explicit visual evidence annotations from widely accepted Grounded VideoQA datasets, including Next-GQA nextgqa , E.T. Bench etbench , and REXTIME rextime . For the specific implementation, we slightly extend ($\sigma$ is set to 20) and manually mask these explicit visual evidence segments in videos.
Next-GQA comprises 8,911 QA pairs and 1,557 videos from NextQA nextqa which focuses on complex temporal and causal reasoning. E.T. Bench includes 10,000 grounded QA samples, requiring models to demonstrate event-level reasoning. REXTIME consists of 12,000 QA samples sourced from ActivityNet acititynet and QV-Highlights qv-highlight and is to evaluate the ability to perform temporal reasoning.
Data Cleaning. To ensure the necessity of explicit visual evidence and the uniqueness of the correct answer ${a}^{\ast}$ for the answer set $\mathbb{A}$, we conduct additional data filtering:
1. We exclude questions that could be correctly answered purely through commonsense reasoning. During commonsense exclusion, only questions $q$ and answer sets $\mathbb{A}$ are assessable. We exclude the examples that can be answered correctly without visual inputs $v$ (e.g., “Q: How does the boy feel after he has gotten up the first time he successfully rolled - A: Happy”). 2. We exclude questions focused on simple description questions and temporal location questions. Descriptive questions refer to questions about non-inferential details, such as the color or quantity of sth. 3. We exclude questions that are insufficient to be reasoned out with contextual visual information. Insufficient answer exclusion refers to questions that cannot be deduced from context, such as those beginning with ‘When’ asking for a time, or ‘Where’ inquiring about a location. Furthermore, we exclude questions that are manually identified as undeducible from context, detailed in the Appendix.Sec. E.
These exclusions ensure that the correct answer $a^{\ast}$ of each question can be uniquely reasoned out, avoiding ambiguity in answers. Based on the three principles, we collect 5,549 QA samples, where 3,749 samples are for the training set and 1,800 samples are split into the test set.
Figure 3: Dataset Cleaning and Annotation Pipeline. We carefully check the rigor of the implicitness and filter context QA pairs with no overlap to mitigate potential information leakage. We further rewrite them as clue pairs and label the relations between acquiring clues and answering the question.
Data Annotation. As discussed above, effectively leveraging clues is crucial when tackling implicit video questions. To support the following research, we have converted the QA pairs toward the same video into context clue annotations. These context QA pairs do not have overlapping durations with the target implicit question, as indicated by the blue arrows in Fig. 3. The non-overlapping condition is imposed to mitigate potential information leakage, ensuring that the clues do not directly contain parts of the answer. Specifically, we rewrite the context question into a declarative sentence as the action, and introduce the correct answer to the question as the intent as analyzed in IntentQA intentqa .
Figure 4: Dataset Statistics. The majority of videos have a duration of around one minute, with explicit evidence typically accounting for 10-30% of the video. Questions starting with ‘Why,’ ‘How,’ and ‘What’ are common. Most of the context clues transformed from context QA pairs are beneficial for answering implicit questions, with the majority containing 3-4 context clues.
Moreover, to enable our work and future research to more effectively utilize these contextual clues, we annotate the relations $r=[r_{1},...,r_{n}]$ between the context pair $1\sim n$ and the implicit question $q$ using GPT-4, following the causal relation definition in REXTIME rextime and MECD MECD . Specifically, we determine whether a causal relation exists between the acquisition of context clues and the deduction of the implicit answer. Additionally, we instruct GPT-4 to pay attention to using the counterfactual inference while avoiding the illusory existence causality MECD . We label context clues contributing to implicit question answering as “1” and those that don’t as “0”. Furthermore, we manually validate the quality of 200 causal relationship annotations and find that they highly align with human expert judgment (major choice of 10 undergraduate students), with a similarity of over 92%.
Data Statistics. Data statistics are shown in sub-figures and the caption of Fig. 4. Most questions are centered around “Why” (e.g., “Why did the man in black in the kitchen start laughing?”), “How” (e.g., “How does the man engage with the women and showcase his culinary skills?”), and “What” (e.g., “What did the ladies do after they passed the men some things?”). Our dataset highlights that each video contains multiple valuable context clues, which can serve as effective reasoning paths.
4 Methods
In this section, we present IRM as shown in Fig. 5, which mainly consists of the visual encoder, AIM (Action-Intent Module) (Sec. 4.1), VEM (Visual Enhanced Module) (Sec. 4.2), and the LLM.
4.1 Action Intent Module
As discussed above, based on the theories from cognitive science, we construct contextual clue information as an additional reasoning pathway to assist implicit reasoning tasks. Specifically, leveraging the powerful basic recognition and reasoning capabilities of IRM (initialized from VideoChat2 videochat2 after vision-language alignment and connection), we prompt it with instruction prompt P (shown in the left-top legend of Fig. 5P: Task: Please extract …) to generate action-intent clue candidates.
Since the visual evidence in the context is explicit, extracting these direct, shallow intentions from the actions is much less challenging than directly extracting the indirect, implicit intentions behind the question. LLM is prompted with the uniformly sampled frames of implicit visual input $v$ and the In-Context Learning prompt. Context action sets $\mathbb{A}_{c}=\{act^{1},...,act^{n}\}$ and corresponding intents set $\mathbb{I}=\{int^{1},...,int^{n}\}$ are generated, which are then embedded and concatenated as follows:
$\small X_{A}=\texttt{Cat}(f_{T}(act^{1}),...,f_{T}(act^{n})),X_{I}=\texttt{Cat}(f_{T}(int^{1}),...,f_{T}(int^{n}))$ (1)
where $f_{T}$ indicates the word embedding layer, Cat indicates concatenate. To further refine the clues, we focus on mitigating the hallucination problem of clues while reasoning out key clues beneficial for implicit reasoning. The current action clues may be influenced by hallucinations, but this can be effectively mitigated through careful verification using visual information MECD ; hallucination ; mmvu . Therefore, we incorporate cross-action-vision attention to query each action candidate with global visual information $Q_{V}$ extracted by QFormer. Specifically, the mechanism is performed by the following equation:
$\small X_{I}^{V}=\text{$\Phi_{h}^{C}$}(X_{A},f_{P}(Q_{V}))$ (2)
where $\Phi_{h}^{C}$ indicates the cross-attention for hallucination mitigating, the former input of $\Phi_{h}^{C}$ indicates Query, while the latter is for Key and Value. $f_{P}$ indicates the visual projection layers. By further verifying clues with global visual information, we enhance the injection of temporal awareness while sequentially identifying contextual clues to mitigate hallucinations.
Moreover, action-intent clue pairs contribute differently to the reasoning of the implicit question. Therefore, we introduce the causal relations classifier with implicit question $q$ to determine which intent-action pairs assist in reasoning. Specifically, we concatenate the visually enhanced $X_{I}^{V}$ and the original intent tensor $X_{I}$ with the embedded question $f_{T}(q)$ and then pass them to the causal attention classifier MECD for relationship deduction. This process is represented as:
$X_{I}^{Q}$ $=\texttt{Cat}(X_{I},f_{T}(q)),X_{I}^{V,Q}=\texttt{Cat}(X_{I}^{V},f_{T}(q))$ (3)
$\hat{r}$ $=g_{r}(\texttt{{Cat}}(\text{$\Phi_{r}^{C}$}(X_{I}^{Q},X_{I}^{V,Q}),\text{$\Phi_{r}^{S}$}(X_{I}^{Q})))$
where $\Phi_{r}^{C}$ and $\Phi_{r}^{S}$ represents cross-attention and self-attention, and $g_{r}$ represents the linear layer. Relational reasoning is achieved through causal attention mechanisms, which model the deep causal relationship between each clue and the question—that is, whether the clue contributes to answering the question. We introduce the relation loss $\mathcal{L}_{R}$ to supervise the reasoning process during training:
$\small\mathcal{L}_{R}=\mathcal{L}_{ce}(r,\hat{r}*r_{m})$ (4)
where $\mathcal{L}_{ce}$ indicates the cross-entropy loss, $r_{m}$ indicates the matching results. The value of $r_{m}$ is 0 when the action candidate does not match any action pairs annotation, or when the action matches while the deduced intent is incorrect according to the intent annotation. Otherwise, the value of $r_{m}$ is 1. The matching or not is based on the automatic annotations from GPT-4. During inference, the refined context action clues $X_{A}^{\prime}$ and intent clues $X_{I}^{\prime}$ can be represented as follows:
$\small X_{A}^{\prime}=\texttt{Cat}\left({X_{A}}[i]\right),X_{I}^{\prime}=\texttt{Cat}\left({X_{I}}[i]\right),~{}\hat{r}[i,0]>\hat{r}[i,1].$ (5)
By retaining only the clues inferred to be relevant to the question, AIM effectively filters out Action-Intent clues that are semantically accurate and conducive to implicit reasoning.
Figure 5: IRM Framework. IRM employs dual-stream modeling of contextual actions and intent clues as implicit reasoning chains. Dual clues are generated and refined by AIM through hallucination mitigation and clue-question relationship deduction. Meanwhile, VEM takes the refined clues to enhance visual input. Finally, LLM is prompted with dual clues to conduct implicit reasoning.
4.2 Visual Enhanced Module
As discussed above, the VEM first introduces QFormer blip2 to compress extensive visual input $f_{V}(v)$ from the visual encoder, represented as $Q_{V}$. During the compression, the instruction to Qformer is: “Extract the context clues related to the question: q”. Following the previous works videochat2 ; videollava ; videollama ; compress , compressed visual input $Q_{V}$ is then sent to visual projection layers $f_{P}$ to align with textual embeddings, represented by $X_{V}=f_{P}(Q_{V})$.
The aligned visual input is then sent to the AIM module for dual context clues reasoning. Meanwhile, the current visual extraction $X_{V}$ only implicitly suggests the need to extract key context visual clues. Therefore, for explicit visual enhancement, we establish a feedback path from AIM to VEM, utilizing context clues generated by AIM to enhance the visual extraction of VEM:
$\small X_{V}^{\prime}=\text{$\Phi_{v}^{C}$}(X_{V},\texttt{Cat}(X_{A}^{\prime},X_{I}^{\prime})).$ (6)
where $\Phi_{v}^{C}$ indicates the cross-attention. Therefore, by further querying the selected key textual clues within the visual input, clue-enhanced visual embeddings are obtained. Additionally, as shown in the simplified diagram in Fig. 5, the coupled design of AIM and VEM enables additional enhancement iterations during inference. Specifically, the global visual input $f_{P}(Q_{V})$ of Eq. 2 can be replaced by enhanced $X_{V}^{\prime}$ after Eq. 6 through an additional iteration. The LLM finally conducts reasoning:
$\small Ans=LLM(X_{A}^{\prime},X_{I}^{\prime},X_{V}^{\prime},q).$ (7)
In this way, the model effectively utilizes both textual and visual clues of context when answering implicit questions. Detailed prompts can be referred to in the Appendix.Sec. D.
5 Experiments
Table 1: Results on I-VQA. IRM reaches SOTA performances in both multi-choice and open-ended settings. Clues are fairly provided by default, * indicates w/o provided.
Paradigm Method Multi-choice Open-ended
Accuracy Score $\text{Acc}_{\text{open}}$
- Random Guess 21.14 - -
Zero-Shot Reasoning LLM
DeepSeek-R1 deepseek 49.75 3.23 54.09
OpenAI-o3 o3 53.77 3.20 55.35
Proprietary ImageLLM
[…]
Gemini-2.5-Pro gemini 52.84 3.25 53.57
Claude-3.7-Sonnet claude 53.31 3.27 52.83
GPT-4o gpt4 54.38 3.28 54.41
Open-Source MLLM
Minigpt4-video Minigpt4-video 35.85 2.88 41.60
VideoLLaVA videollava 36.26 2.94 42.80
PLLaVA pllava 44.38 3.11 47.93
Qwen2.5-VL-7B qwen 46.95 3.13 48.72
NViLA nvila 48.42 2.95 46.08
VideoChat2 videochat2 50.27 3.15 52.24
Fine-tuning Open-Source MLLM
PLLaVA pllava 43.82 3.17 52.46
Qwen2.5-VL-7B qwen 51.26 3.26 54.32
VideoChat2 videochat2 51.90 3.27 53.79
Self-IRM 54.50 3.28 55.31
IRM 55.14 (+4.87) 3.29 56.02
- Human 92.50 - -
Figure 6: Accuracy and Efficiency Trade-off Results. Our model achieves both optimal performance and high efficiency.
Figure 7: Visualizations of multi-choice setting of I-VQA. Results indicate the powerful reasoning ability of IRM.
5.1 Implementation Details
Architectures and Training. We utilize VideoChat2 videochat2 as baseline with UMT-L umt as visual encoder and Mistral-7B mistral as LLM. We initialize IRM after vision-language alignment and connection to conduct instruction tuning. We trained IRM on 8 A6000 GPUs for about 3 hours. We use GPT-4o by default for candidate clue generation during training and IRM for inference, while the Self-IRM indicates using IRM for both training and inference. More details are in the Appendix.Sec. C.
Datasets. We conduct experiments on the I-VQA and the PSAV dataset adsdataset . The PSAV includes 1,002 examples for identifying the best strategy among 12 advertisement persuasion strategies.
Evaluation Metrics. For the open-ended setting of I-VQA, we adopt the “Score” and “$\text{Acc}_{\text{open}}$” metric following previous VideoQA works based on LLMs videollava ; qwen ; pllava ; stllm , where prediction is scored (from 0 to 5) and judged true or false by GPT-3.5. Besides, the high coherence between “Score” and human is proved in the Appendix.Sec. H, with Pearson correlation coefficient pearson reaching a high score of 0.93. We adopt option “Accuracy” for multi-choice setting. For PSAV, we follow the metric in adsdataset : “Accuracy” to evaluate the answer precision. Moreover, since multiple correct answers may exist for PSAV, we introduce “Recall” to evaluate the proportion of the answers covered.
Comparative Methods. We compare IRM with powerful Proprietary ImageLLMs GPT-4o gpt4o , Gemini-2.5-Pro gemini , and Claude-3.7-Sonnet claude . Similarly, we evaluate the leading open-source MLLMs, VideoLLaVA videollava , PLLaVA pllava , Minigpt4-video Minigpt4-video , NViLA nvila , Qwen2.5-VL qwen , and VideoChat2 videochat2 . We also evaluate reasoning models OpenAI-o3 o3 and DeepSeek-R1 deepseek . For fairer comparison, we fine-tune PLLaVA, Qwen2.5-VL, and VideoChat2. To ensure fairness, all comparison models are also prompted with the same context clues generated by GPT-4o during reasoning, which is strictly aligned with our IRM. Fine-tuning details and prompts for IRM and other models are in the Appendix.Sec. D, specific API version details are in the Appendix.Sec. C.
5.2 Main Results
Experiments of Invisible Evidence Questions.
As shown in Tab. 1, our IRM achieves SOTA in both the multi-choice and open-ended settings of the I-VQA dataset, with an accuracy of 55.14 and a GPT score of 3.29. This surpasses existing MLLMs, including fine-tuned VideoChat2 videochat2 , as well as GPT-4o gpt4 . Specifically, IRM demonstrates improvements of over 0.67% in $\text{Acc}_{\text{open}}$ and 0.76% in accuracy compared to the closest OpenAI-o3 and GPT-4o, exhibiting slightly stronger implicit reasoning abilities than strong commercial models. Accuracy and efficiency trade-off is shown in Fig. 5, detailed statistics are in the Appendix.Sec. B.
[…]
Reasoning LLM OpenAI-o3 and DeepSeek-R1 demonstrate strong reasoning abilities under the enhancement of contextual clues generated by GPT-4o, successful and failed examples are in the Appendix.Sec. F. For validating the exclusion of commonsense questions, we also evaluate two reasoning models only with the question and options, the accuracy drops to 23.51 and 22.74, which is only slightly higher than random guessing. This result also indicates a small language bias perception of I-VQA. Visualizations of different settings in Fig. 7 and Fig. 5.2 demonstrate that IRM successfully infers the correct answer with dual refined clues. We also present the subjective results under the open-ended condition for examples that are not included in the open-ended subset. As shown in Fig. 5.2, IRM effectively generates multiple reasonable outcomes along with their probabilities.
Zero-shot Experiments of Implicit Intent Questions.
As shown in Tab. 2, IRM achieves SOTA performance with a recall of 45.41% and an accuracy of 46.08%, significantly outperforming the VideoChat2 by 24.49% and 6.86%, respectively. The results demonstrate IRM’s strong generalization ability and indicating we greatly enhance the reasoning capacity for general implicit questions. Visualizations are in Fig. 10 and the Appendix.Sec. A.
Figure 8: Visualization of the Open-ended Setting of I-VQA. Result indicates the reasoning and clues refining ability of IRM.
Figure 9: Open-ended Result of Example with Non-unique Answers. IRM effectively generates multi answers with probabilities.
Table 2: Results on PSAV Dataset. Our IRM also achieves SOTA performances in the zero-shot test.
Method Recall Accuracy
GPT-3.5 Story + Flan-t5 Classifier adsdataset - 33.41
GPT-3.5 Story + GPT-3.5 Classifier adsdataset - 35.02
OpenAI-o3 o3 36.38 45.84
GPT-4o gpt4 33.51 44.92
DeepSeek-R1 deepseek 23.62 21.96
VideoChat2 videochat2 20.92 39.22
PLLaVA pllava 38.36 39.40
NViLA nvila 38.13 41.09
Self-IRM 39.02 45.14
IRM 45.41 (+24.49) 46.08 (+6.86)
Figure 10: Visualizations of the PSAV Dataset. The results demonstrate the powerful implicit reasoning ability of IRM.
5.3 Ablation Study
Experiments of the Dual Clues and Architecture. To demonstrate the effectiveness of incorporating context clues, we compare: (1) No guidance. (2) Utilizing Grounding-Dino groundingdino for object tracking. (3) Employing PySceneDetect PySceneDetect for scene detection, followed by SceneGraphParser SceneGraphParser to convert the context scenes into a triplet-based graph. As shown in Tab. 3, context dual clues lead to improvements of 3.20%, 2.53%, and 2.36%, respectively. This also proves that the SOTA performance of IRM is not simply achieved through video enhancement methods like partial masking videomae .
Table 3: Ablation Study. We validate the effectiveness of the basic dual clues structure and confirm the refinement and enhancement effect of the proposed coupling AIM and VEM modules.
Context Clues Designs IRM Designs Acc
Clue Tracking Scene Graph Instruction AIM VEM
51.94
52.61
52.78
53.28
53.50
53.67
53.00
55.14
0.20.20.20.20.30.30.30.30.40.40.40.40.50.50.50.50.60.60.60.60.70.70.70.70.80.80.80.80.90.90.90.9111148484848515151515454545457575757Percent (Unit 1 for 3.7k Data / 32 Frames / 5 Iters)Accuracy (%)
Figure 11: Training Data, Reasoning Time and Input Frames Ablation Study. Inference accuracy increases with more training data (red) and more reasoning time (green). Optimal performance is reached with 8 input frames (blue) considering efficiency.
Table 4: Visual Encoder and LLM Ablations. The greater impact of encoder indicates the importance of visual clues extraction.
     Visual Encoder      LLM      Acc
     EVA-CLIP-g eva      Vicuna-7B v1.5 vicuna      51.69
     EVA-CLIP-g eva      Mistral-7B  mistral      51.98
     UMT-L umt      Vicuna-7B v1.5 vicuna      54.31
     UMT-L umt      Mistral-7B mistral      55.14
Figure 12: Visualization of Traffic-VQA. The correct answer, “most likely not” is inferred by our IRM with context clues: the trucks braking to avoid the white.
Table 5: Downstream Task Results of the Traffic-VQA Dataset. IRM’s SOTA performance in zero-shot indicates generalizability.
           Method            Accuracy
           GPT-4o gpt4            45.83
           OpenAI-o3 o3            47.15
           PLLaVA pllava            37.45
           VideoChat2 videochat2            43.76
           DeepSeek-R1 deepseek            45.81
           NViLA nvila            46.12
           Self-IRM            46.22
           IRM            47.94 (+4.18)
Besides, we evaluate our design of VEM for enhancing visual information extraction. We also evaluate the coupled AIM and VEM, while the improvement is modest when introduced individually, combining both modules results in a significant boost.
Experiments of Data, Reasoning Time and Frames. As shown in Fig. 11, accuracy continues to rise while increasing the data volume during SFT, similar to the post-training scaling law aligner ; scaling .
We also study the relationship between reasoning time and accuracy during inference. By iteratively enhancing the visual and textual clues, as the green line in Fig. 11, performance improves with the additional 1-3 iterations (from 55.14 to 55.59, 55.73, 56.04, indicating the inference scaling law with reasoning time time ; scaling . Specifically, increasing iterations results in a better performance compared to traditional reasoning methods such as Greedy Search greedy of 55.65 and Majority Voting vote of 55.87 while maintaining the same FLOPs. Moreover, IRM reaches optimal performance with 8 frames considering efficiency, which may be related to the pre-training process.
Experiments of Visual Encoder and LLM. We default to utilizing UMT-L and Mistral-7B as our visual encoder and LLM, respectively, while also evaluating other models such as EVA-CLIP eva for the encoder and Vicuna vicuna for the LLM. As shown in Tab. 5.3, the clues extraction ability of the visual encoder has a greater impact. This shows that, in implicit reasoning, the context clues, which serve as additional textual information and enable the extraction of visual clues, are more important than the subsequent reasoning.
5.4 Generalization Evaluation
Moreover, we introduce the traffic-VQA for generalization evaluation, where models are required to trace the causes of traffic accidents. Based on the SUTD-Traffic dataset sutd , we select the “Event Forecasting” subset because it evaluates on predicting implicit outcomes. Finally, we collect 2,195 examples for generalization evaluation in a zero-shot manner. As shown in Tab. 5, IRM also performs SOTA, surpassing VideoChat2 and GPT-4o by 4.18% and 2.11%, visualization is shown in Fig. 12.
To further demonstrate the generalizability in explicit questions, we evaluate IRM on four sub-tasks of the widely accepted video understanding benchmark MVbench videochat2 (detailed in the Appendix.Sec. G). The results show that IRM outperforms VideoChat2 by 1.88% on accuracy. In particular, IRM exhibits a significant performance gain in “Action Prediction” (+8.6%) and “Moving Direction” (+4.1%), indicating that implicit video reasoning is inherently connected to predicting unknown future events.
6 Conclusion
We proposed a novel task, Implicit Video Question Answering (I-VQA), which focuses on video reasoning when explicit visual evidence is not directly accessible or locatable. Besides, we built the I-VQA dataset and proposed the first implicit VideoQA framework IRM, which introduces dual context clues as auxiliary reasoning paths. Experimental results demonstrate that IRM outperforms previous SOTA MLLMs, including GPT-4o and VideoChat2. Furthermore, IRM shows strong performance on advertisement understanding and traffic-VQA, highlighting its generalizability across a wide range of video reasoning tasks and its potential for video event prediction.
Appendix
Appendix A More Visualizations
Figure 13: Visualizations of the Multi-choice Setting of the I-VQA Dataset. The intention of the host’s greeting and the action of the woman looking at the cake making “taking a photo” to be most reasonable. In the bottom, as the boy plays with a toy and his little sister shows interest, he will likely give it to her for playing next.
Figure 14: Visualizations of the Open-ended Setting of the I-VQA Dataset. Our IRM integrates contextual information more effectively while reasoning.
In the first visualization example of the multi-choice setting of I-VQA in Fig. 13, the host’s gesture of greeting and hugging the guest indicates the end of an activity, not the beginning, thus ruling out options such as singing, dancing, or musical performance. Furthermore, the unengaged action between the female guest and the cake suggests that the cake hasn’t been cut yet ruling out the intention of eating cake. Based on this reasoning, the most plausible scenario is the appearance of the man in black, who is taking a photo of the host and guest. In the second example, the boy is playing with a toy when his little sister approaches and shows interest in it. Given this interaction, it is highly likely that the boy will give the toy to his sister to play with next.
Figure 15: Visualizations of the SUTD-Traffic Dataset. The routes and intentions of the two vehicles aid in the reasoning of traffic accident prediction.
Figure 16: Visualizations of the PSAV Dataset. Our IRM performs better in both precision and recall when reasoning toward the implicit intent questions.
In the visualization example of the open-ended setting of I-VQA in Fig. 14, the context of many people gathering for a meal creates a joyful atmosphere, suggesting that an implicit action, such as waving to the camera, is likely an indication of the intention to capture a happy moment rather than a request for help. In the SUTD-Traffic dataset experiment shown in Fig. 15, IRM successfully identifies the routes and intentions of the two vehicles, accurately predicting the subsequent unseen traffic accidents.
In the advertisement understanding example shown in Fig. 16, from the perspective of a pair of grandparent and grandchild, the advertisement highlights the social impact of smart agriculture and the public’s recognition of it. Our IRM effectively uncovers all the persuasion strategies underlying the event by analyzing the actions and intentions involved. In contrast, other MLLMs fail to grasp the specific dynamics of the grandparent-grandchild interaction or the promotional aspects of smart agriculture.
Appendix B Inference Speed
As shown in Tab. 6, the computational overhead introduced by our IRM’s context clue refinement and visual clue enhancement is relatively lightweight compared to the overall reasoning process, each sample only takes an additional 0.27 seconds.
[…]
Model Speed (seconds/sample) Model Size
VideoChat2 [46] 1.68 7B
PLLaVA [41] 1.79 7B
VideoLLaVA [39] 2.02 7B
Qwen2.5-VL [40] 2.27 7B
IRM 2.51 7B
NViLA [56] 2.55 8B
IRM with 3 Extra Iterations 3.32 7B
Minigpt4-video [55] 8.94 7B
DeepSeek-R1 [50] ${}~{}\text{9.71}^{*}$ 671B
OpenAI-o3 [51] ${}~{}\text{10.70}^{*}$
[…]
Appendix C More Implementation Details
Architectures and Training. We fine-tune the visual encoder, the QFormer, and the LLM with question generation loss and relation loss, with initial weights of 1.0 and 2.0. The weight of the relation loss decays by 0.05 each epoch, ensuring that the focus of learning gradually shifts to reasoning through clues.
Performances. We report the average results of IRM and all open-source models under three random seeds (2024, 2025, 2026). The default temperature is used for all proprietary models, and the average performance is taken from three experimental runs.
Fine-tuning Details of Comparative Methods. For the fine-tuning of Qwen2.5-VL [40], we followed the official default implementation by freezing the visual encoder to preserve prior knowledge while keeping the LLM parameters learnable. Similarly, for VideoChat2 [46] and PLLaVA [41] fine-tuning, we also followed the official implementations: freezing the visual encoder while training the (Q-Former for VideoChat2) vision-language projection layers, and applying LoRA adapters tuning to the LLM components.
Proprietary Model Version and Input. OpenAI-o3 represents the version of “o3-2025-04-16”, Gemini2.5-Pro represents the version of “gemini-2.5-pro-preview-05-06”, Claude-3.7-Sonnet represents the version of “claude-3-7-sonnet-20250219”, GPT-4o represents the version of “gpt-4o-2024-08-06”. The inputs of Gemini2.5-Pro, GPT-4o, and Claude-3.7-Sonnet are all 8 frames to ensure a fair comparison.
Appendix D Details of Prompt
For the implicit question evaluations on the I-VQA (multi-choice) and the SUTD-Traffic [73] dataset, we design the guiding prompt for IRM and the other comparison models (e.g., GPT-4o) as follows:
"The question involves implicit visual information, with key visual evidencebeing invisible, requiring the deduction of the answer based on contextualvisual information and provided intention, action clues""Context clues are as follows: ""Clues1: XXX to XXX""Clues2: XXX by XXX""Clues3: XXX before XXX""Based on the clues, select the option that accurately addresses the question.""Question: XXX""(A) XXX ... (E) XXX""Only give the best option."
For the open-ended implicit question evaluations on the I-VQA dataset, we design the guiding prompt for IRM and the other comparison models (e.g., GPT-4o) as follows:
"The question involves implicit visual information, with key visual evidencebeing invisible, requiring the deduction of the answer based on contextualvisual information and provided intention, action clues""Context clues are as follows: ""Clues1: XXX to XXX""Clues2: XXX by XXX""Clues3: XXX before XXX""Question: XXX"
Besides, for the implicit question evaluation on the PSAV dataset, we design the guiding prompt for IRM and the other models (e.g., GPT-4o) as follows:
"Please choose one advertising strategy that aligns with the intent of the adv-ertisement:(A) Social Identity(B) Concreteness(C) Anchoring and Comparison(D) Overcoming Reactance(E) Reciprocity(F) Foot-in-the-Door(G) Authority(H) Social Impact(I) Anthropomorphism(J) Scarcity(K) Social Proof(L) Unclear""Context clues are as follows: ""Clues1: XXX to XXX""Clues2: XXX by XXX""Clues3: XXX before XXX""Based on the clues, select the option that accurately addresses the question.""Only give the best option."
Additionally, a complete prompt example of IRM under the multi-choice setting of the I-VQA task:
[…]
The prompt templates we used for IRM, GPT-4o, OpenAI-o3 and other models are shown above. For more detailed code implementation, please refer to the open-source code repository https://anonymous.4open.science/r/Implicit-VQA/.
Appendix E Data Clean Example
Figure 17: Data Cleaning Example. We exclude questions that cannot be reasoned using contextual visual information, those that can be adequately answered through commonsense alone, or questions that are simply toward basic descriptions.
As shown in Fig. 17, we kept the question: “What does the boy in red do after reaching the vase?” because, based on the options, we can infer from the context that, earlier, the two boys were holding paddles and walking towards the vase. Besides, the two boys are then seen moving the charcoal in the vase with sticks. Therefore, the boy approaching the vase is likely trying to add something to it. By analyzing the options, it becomes clear that option C is the only logically correct one that can be deduced.
On the other hand, we excluded the question: “What does the lady do at the end of the video?” because, based on the previous context, we can only infer that the boy threw the paddle into the vase. However, the actions of their mother/sister afterward cannot be uniquely deduced from the context. She could either educate the boy, attempt to remove the paddle, try to extinguish the fire or do something else entirely. Therefore, the answer is not uniquely inferable.
Appendix F Reasoning Examples of DeepSeek-R1
We present the successful reasoning process of DeepSeek-R1 for the question “What does the boy in red do after reaching the vase?” shown in Fig. 17:
Okay, so I’ve got this question here that’s a bit tricky, but I’ll try to work through it step by step. Let me start by understanding the context clues provided.
First, there are three main clues:
1. A boy in red runs around the garden: Play a game for fun. 2. The boy in blue uses a wooden stick to stir the contents in the vase: Make the fire burn more fiercely. 3. A woman in pink walks toward the boys: Supervise play.
Now, the question focuses on what the boy in red does after reaching the vase. So, the boy in red is running around, presumably as part of a game, and when he reaches the vase, he does something specific.
Looking at the options: (A) pick it up (B) put at the side to dry (C) puts his wooden paddle in it (D) get off his horse (E) photos taken
The boy in red is playing a game, so maybe he’s part of a game that involves the vase. The boy in blue is using a wooden stick to stir the vase to make the fire burn more. So, perhaps the vase is part of a fire-related game, like a bonfire or something similar.
If the boy in red runs to the vase, what would he do? Option C is about putting a wooden paddle in it. A paddle is like a stick, so maybe he uses it to stir as well, but the boy in blue is already doing that. Alternatively, maybe the boy in red picks it up (option A) to carry it somewhere, but why would he pick it up? Or maybe he puts it aside to dry (B), but that doesn’t seem related to a game. Option D is about getting off a horse, which doesn’t seem connected. Option E is about taking photos, which also doesn’t fit the context of playing a game.
So, considering all this, the most plausible answer is option C: puts his wooden paddle in it.
Ultimately, through the above analysis, DeepSeek-R1 correctly deduced the boy’s intent in approaching the vase, demonstrating implicit reasoning ability supported by contextual clues. Additionally, as shown in Fig. 18, we also provide a failure case of DeepSeek-R1’s reasoning:
Alright, so I have this question about why the boy was on the floor in the middle of the video.
First, the context clues provided are:
1. A man swings the child: play with the child. 2. The child runs towards the man on the grass: seek affection. 3. The man kneels and opens his arms: invite a hug. 4. The man hugs the child: show affection.
So, the scenario seems to involve interaction between a man and a child, specifically the boy. The actions described are playful and affectionate.
The question is why the boy was on the floor. The options are about different reasons. Let me think about each option in relation to the context clues.
Option A: Wanted to roll on grass. Well, the context mentions the man swinging the child, which is a play activity. Maybe they were outside, but the clues don’t specify grass. So, I’m not sure about this one.
Option B: To dance on the floor. Hmm, if the man is playing with the child, maybe they were dancing together. The man kneeling and hugging could be part of dancing moves. That makes sense because dancing often involves being on the floor or moving around.
Option C: Pick up the bicycle. The context doesn’t mention anything about a bicycle. So, probably not this one.
Option D: Someone pushed him. The clues don’t indicate any negative action like pushing. It’s more about affectionate play.
Option E: Fall down. While playing, a child may not fall with the care of his father.
Now, focusing on option B: dancing on the floor. The clues show that the man is actively engaging with the child, which could include dancing. The actions described, like swinging, running towards, kneeling, and hugging, are all part of playful interaction, which could very well include dancing. So, it makes sense that the boy was on the floor because they were dancing together.
DeepSeek-R1 [50] incorrectly assumes that the presence of a father’s care would prevent falling, instead choosing the option of dancing on the floor, which seems superficially related to the “floor”.
Figure 18: Failure Case Example of DeepSeek-R1.
Appendix G MVbench Subset Evaluation
To demonstrate that IRM still performs well in reasoning for explicit questions, we selected the following tasks from the MVbench [46]:
1) “Action Prediction” (AP) in the Action Reasoning.
2) “Moving Direction” (MD) in the Location Reasoning.
3) “Episodic Reasoning” (ER) in the Cognitive Reasoning.
4) “Object Existence” (OE) in the Object Reasoning.
As shown in Tab. 7, IRM significantly improves action prediction and motion direction judgment sub-tasks. This result is consistent with the previous findings in the evaluation of SUTD-VQA [73] dataset, suggesting that as the model’s understanding of implicit questions improves, its prediction and dynamic judgment capabilities also experience substantial enhancement. Moreover, the overall average accuracy increases by 1.88, demonstrating the strong robustness of our IRM. However, there is a slight decline in episodic reasoning and object existence inference, indicating that in tasks where visual information is particularly explicit, our model may experience some performance degradation. Future work incorporating multi-agents may mitigate these issues and bring about improvements.
Table 7: MVbench Subset Evaluation Results. IRM model brings a general improvement in tasks such as action evolution and behavior prediction in video understanding. Overall, the average accuracy increases by 1.88, demonstrating strong robustness.
Model AP MD ER OE Ave
VideoChat2 [46] 47.5 23.0 40.5 58.0 42.25
IRM 56.1 $\uparrow$ 27.1 $\uparrow$ 37.8 55.5 44.13$\uparrow$
+8.6 +4.1 -2.8 -2.5 +1.88
Appendix H Score Coherence Evaluation
To ensure that the GPT-Score we use for open-ended evaluation is reasonable and consistent with traditional VideoQA, we calculated the Pearson correlation coefficient [59] between the average scores given by five annotators and the scores assigned by GPT-3.5. We found that the coefficient reached as high as 0.93 (with Pearson correlation values ranging from -1 to 1, where values closer to 1 indicate stronger correlation), demonstrating that this metric is also applicable to our proposed implicit-VQA.
Appendix I Details of the Related Datasets
I.1 Details of the GroundedVQA Datasets
Next-GQA comprises 10,531 annotated evidence temporal segments, corresponding to 8,911 QA pairs and 1,557 videos from NextQA [42] which consists of a subset focused on complex temporal and causal reasoning. E.T. Bench includes 10,000 grounded QA samples derived from 7,002 self-collected videos, requiring models to demonstrate event-level comprehension and time-sensitive reasoning. REXTIME consists of 12,000 QA samples with annotated answer temporal locations, sourced from ActivityNet [43] and QV-Highlights [44]. It is specifically designed to evaluate the ability to perform temporal reasoning.
I.2 Details of the PSAV Dataset
The “Persuasion Strategies in Advertisement Videos” dataset [37] (PSAV) contributes an important task on the advertisement story understanding task, namely persuasion strategy identification. For this task, the authors collect 1,002 video advertisements from popular brands available on the web publicly and utilize the persuasion strategy labels defined by [38]. The 12 strategies target persuasion strategy set are: Social Identity, Concreteness, Anchoring and Comparison, Overcoming Reactance, Reciprocity, Foot-in-the-Door, Authority, Social Impact, Anthropomorphism, Scarcity, Social Proof, and Unclear.
I.3 Details of the SUTD-Traffic Datasets
The SUTD-TrafficQA [73] dataset is a large-scale collection of 10,080 videos from diverse real-world traffic scenarios, sourced both online (YouTube, LiveLeak, Twitter, Bili-bili) and offline (handheld cameras and car-mounted recorders) to ensure global diversity. It contains 62,535 human-generated question-answer pairs across six complex reasoning tasks: basic event understanding, event forecasting, reverse reasoning, counterfactual inference, introspection, and attribution. These tasks focus on causal inference, event understanding, and multi-level logical reasoning.
Appendix J Limitations and Future Work
1. We have demonstrated that the thinking patterns related to invisible or deep-level information in implicit questions are beneficial for event/action prediction tasks, as they similarly involve predicting the unknown. In the future, it can also be further explored whether including prediction tasks in the training phase could similarly enhance implicit reasoning abilities.
2. We validated the generalization of our model on the implicit video question answering we proposed, as well as on a broader range of video tasks (including advertisement classification, video event prediction, and explicit video question answering). However, we did not validate it on more other tasks, as they are less related, leaving that exploration for future research.
3. Although our model relies on GPT-4o-generated candidate clues during training, our IRM operates completely independently during inference. Additionally, we have open-sourced the contextual clues to facilitate future research.
4. We only conducted experiments on scaling up the training data and inference time. Due to limited computational resources, the investigation of scaling up the model itself is left for future research.